OpenStudy (anonymous):

find an equation of the tangent line to the graph of f at the given point f(x)=x^2 + 1 at point (2,5)

OpenStudy (anonymous):

find the derivative of f(x)

OpenStudy (anonymous):

Yes, the first thing to do is to find the derivative of f(x). In this case, f'(x) = 2x When x = 2, f'(x) = 4. The slope should be +4, or positive 4. If the slope is positive four and it is going through point (2,5), then y = 4x - 3 This makes sense because if (2,5) is plugged back into the equation, they are equal. Here it is. 5 = 4*(2) - 3 5 = 8 - 3 5 = 5 CHECK The answer is y = 4x - 3.